# Language Model Re-rankers are Steered by Lexical Similarities

### v1

@@ Abstract
Language model (LM) re-rankers are used to refine retrieval results for retrieval-augmented generation (RAG). They are more expensive than lexical matching methods like BM25 but assumed to better process semantic information. To understand whether LM re-rankers always live up to this assumption, we evaluate 6 different LM re-rankers on the NQ, LitQA2 and DRUID datasets. Our results show that LM re-rankers struggle to outperform a simple BM25 re-ranker on DRUID. Leveraging a novel separation metric based on BM25 scores, we explain and identify re-ranker errors stemming from lexical dissimilarities. We also investigate different methods to improve LM re-ranker performance and find these methods mainly useful for NQ. Taken together, our work identifies and explains weaknesses of LM re-rankers and points to the need for more adversarial and realistic datasets for their evaluation.
Language Model Re-rankers are Steered by Lexical Similarities
@@ 1 Introduction
Retrieval-augmented generation (RAG) is used to alleviate problems arising from imperfect parametric knowledge of language models (LMs) (Gao et al., 2024; Vu et al., 2024). However, the efficiency of RAG hinges on the retrieval of useful information (Wang et al., 2024b). To this end, LM re-rankers are increasingly used to provide more accurate retrieval results for RAG, superseding simpler methods based on keyword matching, such as BM25 (see Figure 1). While there are many benchmark results for LM re-rankers (Thakur et al., 2021; Petroni et al., 2021), few extensive inspections of LM re-rankers have been performed. Little is known about when the computationally expensive LM re-rankers are worth the cost and whether they always can be expected to outperform simpler methods.
@@
In this paper, we evaluate LM re-rankers to better understand when they work well and when they fail to outperform less expensive alternatives. The contributions of this paper are as follows:
  • We evaluate 6 LM re-rankers of varying design on the NQ, LitQA2 and DRUID datasets to compare re-ranker performance for scenarios of varying aspects of difficulty and domain.
@@
Taken together, our paper identifies and measures novel aspects of difficulty for LM re-rankers; distractors and lack of contextual information. These aspects are likely to occur in real-world scenarios relying on e.g. information retrieval from the Internet. Our work points to the need of more adversarial and real-world aligned evaluation datasets to better understand and address LM re-ranker fallacies related to the identified aspects of difficulty.
@@ 2 Related Work
In contrast to these works, we expand on the analysis of distractors and missing document context to include multiple SOTA re-rankers, datasets from diverse domains and better tuned metrics. We also tie these aspects of difficulty to a more fundamental question of whether LM re-rankers are steered by lexical similarities. To measure this, we develop a new metric which allows us to identify problematic samples.
@@ 3.1 Re-rankers
We evaluate a wide cohort of LM re-rankers to enable comprehensive comparisons between different model types and sizes. Three closed-source LM re-rankers are evaluated: The industrial grade re-ranker Cohere111rerank-english-v3.0 from https://docs.cohere.com/v2/docs/models#rerank (Cohere), the LLM-based re-ranker GPT-4o (GPT-4o) and the lightweight LLM re-ranker GPT-4o mini (GPT-4o m) (Sun et al., 2023) (Appendix E).222gpt-4o-2024-08-06 and gpt-4o-mini-2024-07-18.
We also evaluate three open-source re-rankers from Hugging Face: The large-scale LM re-ranker bge-reranker-v2-gemma (BGE), the lightweight re-ranker jina-reranker-v1-turbo-en (Jina turbo) and jina-reranker-v2-base-multilingual (Jina base), a larger re-ranker from the same model family. Our baseline is a re-ranker based on BM25 scores that leverages lexical matching, similar to TF-IDF (Lù, 2024). See Appendix D to get a rough estimate of the runtime of each re-ranker.
@@ 3.2 Evaluation datasets
We evaluate the re-rankers on three datasets representative of different domains and aspects of difficulty: NQ, LitQA2 and DRUID. Natural Questions (NQ) is a popular dataset for re-ranker evaluations with passages from Wikipedia pages (Kwiatkowski et al., 2019). LitQA2 measures the ability of a system to extract information from scientific literature (Laurent et al., 2024). The dataset contains a high rate of domain-specific biomedical language and can be expected to test the robustness to domain-shifts of LM re-rankers. DRUID (Dataset of Retrieved Unreliable, Insufficient and Difficult-to-understand contexts) contains fact-checked claims and corresponding potential evidence automatically retrieved from the Internet (Hagström et al., 2024). It can be expected to contain more noisy passages and to test the capability of re-rankers to identify relevant information for fact-checking. More details and examples can be found in Appendix C.
@@ 3.3 Evaluation metrics
We mainly use Precision@1 ($\mathrm{P@1}$) for our re-ranker evaluations to accommodate the small number of passages available in DRUID. [^3] Metrics are defined by TREC in https://trec.nist.gov/pubs/trec16/appendices/measures.pdf. To understand when LM re-rankers fail to outperform simpler methods, we also compare to alignment with BM25 relevance scores. This is measured as follows.
@@
Given re-ranker predictions $R$, $\mathrm{P@1}(R)$ denotes the score measured when document relevance is given by gold labels (default) and $\mathrm{P@1}_{\mathrm{BM25}}(R)$ when relevance is given by BM25 scores. Leveraging this metric, we can investigate whether re-rankers align with gold labels over BM25 scores, which corresponds to positive $\Delta\mathrm{P@1}$ values and vice versa for negative values.
@@ 3.5 Alleviation methods
We investigate two known methods previously shown to improve re-ranker performance: prepending page titles (Prepend titles) (Wang et al., 2024a) and incorporating contextual information generated by GPT-4o mini (Incorporate context). [^4] https://www.anthropic.com/news/contextual-retrieval Prepending titles is quite straightforward for NQ and LitQA2, while the more noisy webpage text in DRUID yields low-quality titles, with missing values and inaccuracies. We were also unable to obtain contextual information for the DRUID passages due to computational limitations. Lastly, we also experiment with adjusting the re-ranker prompt to better suit the fact-checking setting represented by DRUID (Prompt) (Appendix F).
@@ 4 Results
The best-performing re-ranker (BGE) outperforms the BM25 re-ranker by 10% on DRUID. This is smaller than the 46% on NQ (for GPT-4o) and 15% on LitQA2 (for BGE). We also note that the smaller Jina LM re-rankers clearly outperform the BM25 re-ranker on NQ while they perform worse than or equal to BM25 on LitQA2 and DRUID.
@@
We note that while DRUID is easier compared to NQ with respect to LM re-ranker accuracy, it is harder with respect to how LM re-rankers struggle to outperform simpler methods like BM25. We hypothesise that DRUID provides a greater challenge in this sense as it contains passages from the Internet and popular claims that may have seen frequent discussion, increasing the rate of distractors.
@@ 4.2 Alleviation methods
Table 1 report the results from the investigations described in Section 3.5. We reach the following conclusions.
@@
Prepending page titles to the passages yields performance improvements for large LM re-rankers on NQ and unchanged performance on LitQA2 and DRUID. For LitQA2, this could be caused by the more distracting details from the scientific paper titles (Wang et al., 2024a). For DRUID it likely stems from the noisy webpage titles. Seemingly, the method of prepending page titles is more suitable for nicely formatted datasets. We also observe that the method of incorporating contexts is inferior to prepending page titles.
@@ 5 Conclusion
Our paper identifies and explores an important weakness of LM re-rankers: They struggle to identify gold samples with markedly low BM25 scores. We hypothesise that real-world datasets like DRUID, with passages from the Internet, contain more distractors, resulting in gold samples with low BM25 scores. However, most current datasets for re-ranker evaluation fail to capture this aspect of difficulty and methods for improving LM re-ranker performance are less effective for the noisier LitQA2 and DRUID samples. Our work points to the need of more adversarial and real-world aligned datasets to better understand LM re-rankers and their weaknesses in realistic settings.
@@ Acknowledgments
This work was supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Alvis partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ C.3 DRUID
DRUID (Dataset of Retrieved Unreliable, Insufficient and Difficult-to-understand contexts) by Hagström et al. (2024) contains fact-checked claims and corresponding potential evidence pieces retrieved from the Internet. Each evidence piece has been annotated for whether it contains sufficient information to conclude whether the corresponding claim is true or false. The claims from the dataset are used as questions to the re-rankers and the collected DRUID passages corresponding to the given claim make out the passages for the query. Passages with sufficient information to reach a fact-check verdict, i.e. marked as ‘refuting’ or ‘supporting’, are considered gold and each sample corresponds to at least two potential passages from different webpages, of which at least one has to be gold and at least one not gold. The Cohere re-ranker was used for the automated retrieval of evidence pieces so the samples in DRUID can be expected to be more adversarial in the sense that they already have been pre-selected by a LM re-ranker (and then manually annotated for quality).
@@ Appendix F Adjusted prompt for DRUID
Here, “<claim>” and “<claimant>” are replaced by the corresponding values in DRUID. The results for these prompts can be found in Table 1 and Figure 4.

### v2

# Language Model Re-rankers are Fooled by Lexical Similarities

@@ Abstract
Language model (LM) re-rankers are used to refine retrieval results for retrieval-augmented generation (RAG). They are more expensive than lexical matching methods like BM25 but assumed to better process semantic information and the relations between the query and the retrieved answers. To understand whether LM re-rankers always live up to this assumption, we evaluate 6 different LM re-rankers on the NQ, LitQA2 and DRUID datasets. Our results show that LM re-rankers struggle to outperform a simple BM25 baseline on DRUID. Leveraging a novel separation metric based on BM25 scores, we explain and identify re-ranker errors stemming from lexical dissimilarities. We also investigate different methods to improve LM re-ranker performance and find these methods mainly useful for the more popular NQ dataset. Taken together, our work identifies and explains weaknesses of LM re-rankers and points to the need for more adversarial and realistic datasets for their evaluation.
Language Model Re-rankers are Fooled by Lexical Similarities
@@ 1 Introduction
Retrieval-augmented generation (RAG) is used to alleviate problems arising from imperfect parametric knowledge of language models (LMs) (Gao et al., 2024; Vu et al., 2024). However, the efficiency of RAG hinges on the retrieval of useful information (Wang et al., 2024b). To this end, LM re-rankers are increasingly used to provide more accurate retrieval results for RAG, superseding simpler methods based on keyword matching, such as BM25 (see Figure 1). While there are many benchmark results for LM re-rankers (Thakur et al., 2021; Petroni et al., 2021), little is known about when the computationally expensive LM re-rankers are worth the cost and whether they always can be expected to outperform simpler methods.
@@
In this paper, we evaluate LM re-rankers to better understand when they work well and when they fail to outperform less expensive alternatives. [^1] Our code is available at https://github.com/lovhag/rerankers-and-lexical-similarities. The contributions of this paper are as follows:
  • We evaluate 6 LM re-rankers on the NQ, LitQA2 and DRUID datasets to compare re-ranker performance for scenarios of varying aspects of difficulty and domain. NQ is focused on generic QA, LitQA2 on scientific information extraction and DRUID on claim verification.
@@
Taken together, our paper identifies and measures novel aspects of difficulty for LM re-rankers; distractors and lack of contextual information. These aspects are likely to occur in real-world scenarios relying on e.g. information retrieval from the web, such as in a fact-checking setting. Our work points to the need of more adversarial and real-world aligned evaluation datasets to better understand and address LM re-ranker deficiencies.
@@ 2 Related Work
In contrast to these works, we expand on the analysis of distractors and missing document context to include multiple SOTA re-rankers, datasets from diverse domains and better tuned metrics. We also tie these aspects of difficulty to a more fundamental question of whether LM re-rankers are fooled by lexical similarities. To measure this, we develop a new metric which allows us to identify problematic samples.
@@ 3.1 Re-rankers
We evaluate a wide cohort of LM re-rankers to enable comprehensive comparisons between different model types and sizes. Three closed-source LM re-rankers are evaluated: the industrial grade re-ranker Cohere222rerank-english-v3.0 from https://docs.cohere.com/v2/docs/models#rerank (Cohere), a re-ranker based on GPT-4o (GPT-4o) and a lightweight re-ranker based on GPT-4o mini (GPT-4o m) (Sun et al., 2023) (Appendix E).333gpt-4o-2024-08-06 and gpt-4o-mini-2024-07-18.
We also evaluate three open-source re-rankers from Hugging Face: the large-scale LM re-ranker bge-reranker-v2-gemma (BGE), the lightweight re-ranker jina-reranker-v1-turbo-en (Jina turbo) and jina-reranker-v2-base-multilingual (Jina base), a larger re-ranker from the same model family. As a baseline, we consider BM25 scores, leveraging lexical matching, similar to TF-IDF (Lù, 2024). See Appendix D to get a rough estimate of the runtime of each re-ranker.
@@ 3.2 Evaluation datasets
We evaluate the re-rankers on three datasets representative of different domains and aspects of difficulty: NQ, LitQA2 and DRUID. All datasets have already undergone an initial retrieval step and are thus suitable for the evaluation of re-rankers. Natural Questions (NQ) is a popular dataset for re-ranker evaluations with passages from Wikipedia pages (Kwiatkowski et al., 2019). LitQA2 measures the ability of a system to extract information from scientific literature (Laurent et al., 2024), containing a high rate of domain-specific biomedical language. LitQA2 can be expected to test the robustness to domain-shifts of LM re-rankers. DRUID (Dataset of Retrieved Unreliable, Insufficient and Difficult-to-understand contexts) contains fact-checked claims and corresponding potential evidence automatically retrieved from the web (Hagström et al., 2024). It can be expected to contain more noisy passages and to test the capability of re-rankers to identify relevant information for fact-checking. More details and examples can be found in Appendix C.
@@ 3.3 Evaluation metrics
We mainly use Precision@1 ($\mathrm{P@1}$) for our re-ranker evaluations to accommodate the small number of passages available in DRUID. [^4] Metrics are defined by TREC in https://trec.nist.gov/pubs/trec16/appendices/measures.pdf. To understand when LM re-rankers fail to outperform simpler methods, we also compare to alignment with BM25 relevance scores, as follows.
@@
Given re-ranker predictions $R$, $\mathrm{P@1}(R)$ denotes the score measured when document relevance is given by gold labels (default) and $\mathrm{P@1}_{\mathrm{BM25}}(R)$ when relevance is given by BM25 scores. Leveraging this metric, we can investigate whether re-rankers align with gold labels over BM25 scores, which corresponds to positive $\Delta\mathrm{P@1}$ values. Negative $\Delta\mathrm{P@1}$ values correspond to when the re-ranker predictions align more with BM25 scores over gold labels.
@@ 3.5 Alleviation methods
We investigate two known methods previously shown to improve re-ranker performance: prepending page titles (Prepend titles) (Wang et al., 2024a) and incorporating contextual information generated by GPT-4o mini (Incorporate context). [^5] https://www.anthropic.com/news/contextual-retrieval Prepending titles is quite straightforward for NQ and LitQA2, while the more noisy webpage text in DRUID yields low-quality titles, with missing values and inaccuracies. The DRUID samples also lack complete contexts, barring the Incorporate context method. We instead experiment with adjusting the re-ranker prompt to better suit the fact-checking setting represented by DRUID (Prompt) (Appendix F).
@@ 4 Results
The best-performing re-ranker (BGE) outperforms the BM25 baseline by 10% on DRUID. This is smaller than the 46% on NQ (for GPT-4o) and 15% on LitQA2 (for BGE). We also note that the smaller Jina LM re-rankers clearly outperform the BM25 baseline on NQ while they perform worse than or equal to BM25 on LitQA2 and DRUID.
@@
We note that while DRUID is easier compared to NQ with respect to LM re-ranker accuracy, it is harder with respect to how LM re-rankers struggle to outperform simpler methods like BM25. We hypothesise that DRUID provides a greater challenge in this sense as it contains passages from the web and popular claims that may have seen frequent discussion, increasing the rate of distractors.
@@ 4.2 Alleviation methods
The results from the investigations described in Section 3.5 can be found in Table 1. We reach the following conclusions.
@@
Prepending page titles to the passages yields performance improvements for large LM re-rankers on NQ and unchanged performance on LitQA2 and DRUID. For LitQA2, this could be caused by the more distracting details from the scientific paper titles (Wang et al., 2024a). For DRUID it likely stems from the noisy webpage titles. Seemingly, the method of prepending page titles is more suitable for nicely formatted datasets, such as NQ. We also observe that the method of incorporating contexts is inferior to prepending page titles.
@@ 5 Conclusion
Our paper identifies and explores an important weakness of LM re-rankers: they struggle to identify gold samples with markedly low BM25 scores. We hypothesise that real-world datasets like DRUID, with passages from the web, contain more distractors, resulting in gold samples with low BM25 scores. However, most current datasets for re-ranker evaluation fail to capture this aspect of difficulty and methods for improving LM re-ranker performance are less effective for the noisier LitQA2 and DRUID samples. Our work points to the need of more adversarial and real-world aligned datasets to better understand LM re-rankers and their weaknesses in realistic settings.
@@ Acknowledgments
This work was supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Alvis partially funded by the Swedish Research Council through grant agreement no. 2022-06725. The work was also supported by compute credits from a Cohere For AI Research Grant.
@@ C.3 DRUID
DRUID (Dataset of Retrieved Unreliable, Insufficient and Difficult-to-understand contexts) by Hagström et al. (2024) contains fact-checked claims and corresponding potential evidence pieces retrieved from the web. Each evidence piece has been annotated for whether it contains sufficient information to conclude whether the corresponding claim is true or false. The claims from the dataset are used as questions to the re-rankers and the collected DRUID passages corresponding to the given claim make out the passages for the query. Passages with sufficient information to reach a fact-check verdict, i.e. marked as ‘refuting’ or ‘supporting’, are considered gold and each sample corresponds to at least two potential passages from different webpages, of which at least one has to be gold and at least one not gold. The Cohere re-ranker was used for the automated retrieval of evidence pieces so the samples in DRUID can be expected to be more adversarial in the sense that they already have been pre-selected by a LM re-ranker (and then manually annotated for quality).
@@ Appendix F Adjusted prompt for DRUID
Here, “<claim>” and “<claimant>” are replaced by the corresponding values in DRUID. This prompt is tuned to adapt re-rankers to the fact-checking setting, as opposed to a QA setting. The results for these prompts can be found in Table 1 and Figure 4.